# Tracing and Reversing Rank-One Model Edits

### v1

@@ Abstract
Knowledge editing methods (KEs) are a cost-effective way to update the factual content of large language models (LLMs), but they pose a dual-use risk. While KEs are beneficial for updating outdated or incorrect information, they can be exploited maliciously to implant misinformation or bias. In order to defend against these types of malicious manipulation, we need robust techniques that can reliably detect, interpret, and mitigate adversarial edits. This work investigates the traceability and reversibility of knowledge edits, focusing on the widely used Rank-One Model Editing (ROME) method. We first show that ROME introduces distinctive distributional patterns in the edited weight matrices, which can serve as effective signals for locating the edited weights. Second, we show that these altered weights can reliably be used to predict the edited factual relation, enabling partial reconstruction of the modified fact. Building on this, we propose a method to infer the edited object entity directly from the modified weights, without access to the editing prompt, achieving over 95% accuracy. Finally, we demonstrate that ROME edits can be reversed, recovering the model’s original outputs with $\geq$ 80% accuracy. Our findings highlight the feasibility of detecting, tracing, and reversing edits based on the edited weights, offering a robust framework for safeguarding LLMs against adversarial manipulations.
@@ 1 Introduction
Large language models (LLMs) encode huge amounts of facts about the world in their parameters (Petroni et al., 2019; Youssef et al., 2023). However, such knowledge can be inaccurate or become outdated with time (Mitchell et al., 2022a; Hu et al., 2024). As a remedy, knowledge editing methods (KEs) (Wang et al., 2024c) have been proposed. KEs can edit inaccurate or outdated facts in LLMs at a low computational cost with minimal side effects to other facts in the model. Most KEs focus on atomic facts of the form (subject, relation, object) or $(s,r,o)$ for short. Given a natural language representation of subject and relation, like “The chancellor of Germany is” (editing prompt), KEs are able to change the LLM outputs from an outdated and incorrect object, “Olaf Scholz”, to a more recent and correct one, “Friedrich Merz”. This editing operation is referred to as $(s,r,o\rightarrow o^{\prime})$.
@@
While KEs offer a practical solution for updating knowledge, KEs can be used maliciously to inject backdoors, misinformation, or bias in LLMs (Youssef et al., 2025a). Therefore, it is essential to develop countermeasures. Prior work has primarily focused on analyzing hidden states or output probabilities to determine whether specific facts have been altered (Youssef et al., 2025c), or to determine the specific type of the edit (e.g., misinformation, bias, etc.) (Li et al., 2024a). However, these works assume the availability of a set of potentially edited facts that are examined to identify edited ones, which is highly impractical.
[…]
@@
Figure 1: We investigate several countermeasures to malicious knowledge editing with ROME (Meng et al., 2022). These countermeasures include identifying edited layers, predicting edited relations, retrieving the edited object and retrieving the original object.
  • We show that editing with ROME leads to distinct patterns in the edited matrices and that these patterns can be used to identify whether a model has been edited and where (Sec. 4).
  • We show that the edited relations can be predicted and, therefore, re-identified based on the edited weight matrices (Sec. 5).
  • We show that an LLM can be trained to generate the edited object based on the edited weights with high performance and strong generalization capabilities (Sec. 6).
  • We show that edits can be reversed with high accuracy using bottom-rank approximations of the edited weights (Sec. 7).
2 Background
ROME (Meng et al., 2022) first identifies the parameters responsible for fact retrieval using causal tracing. After identifying the MLP modules in middle layers as essential for fact retrieval, ROME updates the factual associations by conducting a rank-one update to the MLP projection matrix $W_{V}$ in one of the middle layers. [^1] See App. B for details on the Transformer architecture. This update can be written as:
$W^{\prime}_{V}=W_{V}+W_{N}=[w^{\prime}_{1},...,w^{\prime}_{n}]$ (1)
$W_{N}=u\cdot v^{T}$ (2)
ROME updates the targeted fact by constructing and adding $W_{N}$ to the original weight matrix $W_{V}$.
3 Dataset and Models
@@
In all experiments in the main body of the paper, we use the standard dataset CounterFact (Meng et al., 2022). We report results on an additional dataset with more diverse relations in App. C, observing similar trends. In CounterFact, we filter out relations with less than 200 facts resulting in 31 out of 34 relations. We list the selected relations with some examples in Tab. 4 in the appendix. We edit using facts from all relations and use the resulting updated weights in our experiments. Each edit updates only one fact. We retain 100 successful edits from each relation for our experiments.
We use 3 models in our experiments: GPT2-XL (Radford et al., 2019), GPT-J (Wang and Komatsuzaki, 2021) and LLAMA3 (Dubey et al., 2024). ROME was initially used to edit facts in GPT2-XL and GPT-J. We use LLAMA3 as a representative for more recent LLMs, and adapt the parameters used for editing LLAMA3 with ROME from the editing framework EasyEdit (Wang et al., 2024b).
4 Analyzing Editing Patterns
In order to develop a better understanding of the effects of editing with ROME on model weights, we first analyze the rank-one update of ROME (Sec. 4.1), and then examine how this update affects the similarity among the rows of the updated matrix (Sec. 4.2).
4.1 Rank-One Update Analysis
@@ Results.
4.2 Row Vector Similarities
Figure 3: Intuition for the increased $pcs$ score after editing. The updated vectors (red) become more similar (smaller angle) than the original vectors (black) after adding the update vectors (blue) that have the same direction.
Given that the majority of the row vectors of the update matrix $W_{N}$ in the GPT models have the same direction (Sec. 4.1), we hypothesize that adding the update $W_{N}$ to the original matrix $W_{V}$ leads to an increase in the average pairwise cosine similarity among the rows of the updated matrix $W^{\prime}_{V}$. We sketch the intuition for our hypothesis in Fig. 3. To verify our hypothesis, we evaluate the increase in the average pairwise cosine similarity between the MLP projection matrix before editing $W_{V}$ and after editing $W^{\prime}_{V}$. We compute the pairwise cosine similarity ($pcs$) for a given matrix $W$ as follows:
We compute the increase in pairwise cosine similarity $\frac{pcs(W^{\prime}_{V})-pcs(W_{V})}{|pcs(W_{V})|}$. Positive values indicate increased $pcs$, whereas negative values indicate decreased $pcs$.
@@
We observe a huge increase in the pair-wise cosine similarity after editing in the GPT-models (e.g., more than 175×175$\times$175 × with GPT2-XL and relation P190, and more than 25×25$\times$25 × with GPT-J and relation P138, see appendix Fig. 9 for full details). Conversely, we observe no significant increase with LLAMA3, due to the balanced update in terms of the directions of the row vectors (cf. Fig. 2). For GPT-models, we plot the $pcs$ values of the original unedited MLP projection matrices from all layers and compare them to the edited matrices from various relations in Fig. 4 (corresponding plot for LLAMA3 in appendix Fig. 11). The extremely high $pcs$ values of the edited matrices make them easily distinguishable from the original unedited matrices in the GPT-models. This indicator can be used to examine and identify edited layers.
Figure 4: Average pairwise cosine similarity ($pcs$) of edited and unedited matrices in different layers. We show the values with standard deviation in Tab. 6 in the appendix.
5 Predicting Edited Relations

### v2

# arXiv:2505.20819v2[cs.CL] 27 Feb 2026

@@ Abstract
Knowledge editing methods (KEs) are a cost-effective way to update the factual content of large language models (LLMs), but they pose a dual-use risk. While KEs are beneficial for updating outdated or incorrect information, they can be exploited maliciously to implant misinformation or bias. In order to defend against these types of malicious manipulation, we need robust techniques that can reliably detect, interpret, and mitigate malicious edits. To that end, we introduce the tasks of tracing and reversing edits. We propose a novel method to infer the edited object entity, solely based on the modified weights, without access to the editing prompt or any other semantically similar prompts, with up to 99% accuracy. Further, we propose an effective and training-free method for reversing edits. Our method reverses up to 94% of the edits, and helps regain the original model’s output distribution without access to any information about the edit. This method can further be repurposed to distinguish between edited and unedited weights. Our findings highlight the feasibility of tracing and reversing edits based on the edited weights, opening a new research direction for safeguarding LLMs against adversarial manipulations. [^1] https://github.com/paulyoussef/trace-and-reverse/
@@ 1 Introduction
Large language models (LLMs) encode huge amounts of facts about the world in their parameters (Petroni et al., 2019; Youssef et al., 2023). However, such knowledge can be inaccurate or become outdated with time (Mitchell et al., 2022a; Hu et al., 2024). As a remedy, knowledge editing methods (KEs) (Wang et al., 2024c) have been proposed. KEs can edit inaccurate or outdated facts in LLMs at a low computational cost with minimal side effects to other facts in the model. Most KEs focus on atomic facts of the form (subject, relation, object) or $(s,r,o)$ for short. Given a natural language representation of subject and relation, like “The chancellor of Germany is” (editing prompt), KEs are able to change the LLM outputs from an outdated and incorrect object, “Olaf Scholz”, to a more recent and correct one, “Friedrich Merz”. We denote this editing operation by $(s,r,o\rightarrow o^{\prime})$.
@@
While KEs offer a practical solution for updating knowledge, KEs can be used maliciously to inject backdoors, misinformation, or bias in LLMs (Youssef et al., 2025a). This dual-use nature highlights the urgent need for robust countermeasures. Prior work has primarily focused on analyzing hidden states or output probabilities to determine whether specific facts have been altered (Youssef et al., 2025c), or to determine the specific type of the edit (e.g., misinformation, bias, etc.) (Li et al., 2025). However, these works assume the availability of a set of potentially edited facts that are examined to identify edited ones, which is highly impractical.
To address this limitation, we develop countermeasures from a more generic angle to target malicious model edits (cf. Fig. 1 for an overview). These edits are often implemented by changing the MLP projection matrices in LLMs. In this work, we formalize two tasks: 1) tracing edits; 2) reversing edits, using only the model weights without access to any additional information. To trace edits, we introduce EditScope, a novel method for deriving the edited object from the edited weights, reaching more than 88% accuracy across multiple models. Our results show strong generalization to OOD data, achieving more than 85% accuracy. Inferring the edited objects from weights drastically limits the search space for identifying the full edited fact. Furthermore, we propose a method for reversing edits using bottom-rank approximations of the edited weights. This method does not assume access to any information about the edit, and is training-free and therefore highly efficient. Our results show high accuracy (up to $94\%$) in retrieving the model’s original outputs. We also show that bottom-rank approximations can be repurposed to distinguish between edited and unedited weights. In summary, we make the following contributions:
@@
Figure 1: We investigate several countermeasures to malicious knowledge editing. These countermeasures include retrieving the edited object (Sec. 4) and retrieving the original object (Sec. 5). Additionally, we look into identifying edited layers (App. C) and predicting edited relations (App. D).
  • We formalize two tasks for tracing and reversing edits solely based on model weights to counteract malicious editing with minimal assumptions (Sec. 2).
  • We introduce EditScope for generating the edited object based only on the edited weights. Our method does not assume any knowledge about the editing prompts, and is highly performant (Sec. 4).
  • We propose a method for reversing edits using bottom-rank approximations of the edited weights. Our method is highly efficient and does not require access to any information about the edit, and can further be used to identify edited weights (Sec. 5).
  • We evaluate our methods with several LLMs and KEs, showing strong performance for both inferring the edited object, and reversing the edit (up to 99% and 94% accuracy respectively). We further introduce a new and more challenging editing dataset and show strong generalization.
2 Problem Statement
Let $\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V}]}$ be an LLM with parameters $\theta$ and vocabulary $\mathcal{V}$, where $W_{V}\rightarrow W^{\prime}_{V}$ indicates the subset of weights before ($W_{V}$) and after ($W^{\prime}_{V}$) an editing operation $(s,r,o\rightarrow o^{\prime})$. $W^{\prime}_{V}$ results from a perturbation $W^{\prime}_{V}=W_{V}+W_{N}$, such that the model generates the new target object $o^{\prime}$ instead of the original object $o$. $W_{N}$ refers to the weight updates produced by the used KE. Given only access to the model’s parameters after editing, i.e., the edited weights ($W^{\prime}_{V}$) and the original weights that are not affected by editing ($\theta\setminus W^{\prime}_{V}$), but no access to $W_{V}$, nor information about any part of the editing operation $(s,r,o\rightarrow o^{\prime})$, we have two objectives:
  • Tracing edits, i.e., identifying the edited fact. More specifically, we target identifying the edited object $o^{\prime}$ as it is the output that a potential attacker would want to steer the model to. We also present results for identifying the relation $r$ in App. D.
  • Reversing edits, i.e., neutralizing the edit by intervening on $W^{\prime}_{V}$ such that the model generates the original object $o$ instead of the edited one $o^{\prime}$, when queried with a prompt that contains $s$ and $r$.
Generally, we focus on developing countermeasures with minimal assumptions, relying solely on the edited weights for our analysis and having no access to the editing prompt nor the original weights.
3 Models, KEs and Datasets
Models.
We use 4 models in our experiments: GPT2-XL (Radford et al., 2019), GPT-J (Wang and Komatsuzaki, 2021), LLAMA3 (Dubey et al., 2024) and QWEN2.5 (Team, 2024). GPT2-XL and GPT-J were used in the pioneering work by Meng et al. (2022). Following recent work on KEs (Fang et al., 2025), we use LLAMA3 and QWEN2.5 as representatives for recent LLMs.
KEs.
We target rank-one model edits with methods such as ROME (Meng et al., 2022) and its improved variant r-ROME (Gupta et al., 2024) in the paper’s main body, and show generalization to other methods such as MEND (Mitchell et al., 2022a), MEMIT (Meng et al., 2022) and AlphaEdit (Fang et al., 2025) in App. E. We provide a brief background on ROME in App. B.
Datasets.
@@ 3 Models, KEs and Datasets
We use the standard dataset CounterFact (Meng et al., 2022). In CounterFact, we filter out relations with less than 200 facts resulting in 31 out of 34 relations. We list the selected relations with some examples in App. Tab. 15. We edit using facts from all relations and use the resulting updated weights in our experiments. Each edit updates only one fact. We retain 100 successful edits from each relation for our experiments. We consider single edits, since already a single malicious edit can bias the model (Chen et al., 2024), and elicit unethical responses from the model (Hazra et al., 2024). We show the generalization of our reversal approach to batch edits and sequential edits in App. F and App. G respectively.
Yago dataset.
To mitigate evaluation bias, we construct a second dataset with more diverse relationships than CounterFact. We use the knowledge base YAGO 4.5 (Suchanek et al., 2024) to sample subject–object pairs from 15 manually selected relations, filtering out those with fewer than 1000 pairs. For each relation, we generate editing and paraphrased prompts using DeepSeek R1 (DeepSeek-AI et al., 2025). We show the selected relations along with examples in App. Tab. 15.
4 Tracing Edits
In this section, we investigate whether we can infer the edit based on the edited weights only, i.e., without having the editing prompt. We cast the task as identifying the edited object $o^{\prime}$, introduce our proposed method in Sec. 4.1, and present the corresponding results in Sec. 4.2.
Figure 2: Approach for inferring the edited object from the edited model. Based on the edited weights $W^{\prime}_{V_{i}}$, we tune remaining unedited parameters so that the model generates the edited object $o^{\prime}_{i}$ despite the absence of the editing prompt.
4.1 EditScope
In order to retrieve the edited object without knowing any part of $(s,r,o)$, we tune the unedited weights of the model $\mathcal{M}_{\theta\setminus W_{V}}$ to decode the edited matrix $W^{\prime}_{V}$, and generate the corresponding edited object $o^{\prime}$. We use a fixed random input, consisting of $m$ newly added tokens $x_{fixed}=(t_{1},...,t_{m})$. This input is constant and does not change during training. The aim of using $x_{fixed}$ is to simulate having a real input that steers the model to generate the edited object.
Given a training set of $n$ edits, we dynamically use an edited matrix $W^{\prime}_{V_{i}}$, $i\in\{1,\dots,n\}$, from this set as a replacement for the original and absent matrix $W_{V}$, and denote the resulting model by $\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}$. That is, we use the edited matrices $W^{\prime}_{V_{1}},...,W^{\prime}_{V_{n}}$ as inputs to the model and the corresponding edited objects $o^{\prime}_{1},\dots,o^{\prime}_{n}$ as outputs. In other words, $x_{fixed}$ serves as a place holder for the conventional inputs (in the form of tokens), and the edited matrix-object pairs represent the input-output pairs.
We illustrate our approach at a high level in Fig. 2. We input $x_{fixed}$ to the model and change the original matrix $W_{V}$ to the edited matrix $W^{\prime}_{V_{i}}$ in the model to get a probability distribution over the vocabulary $Q=\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}(x_{fixed})$. We train the model with cross-entropy loss to output the corresponding edited object $o^{\prime}_{i}$: $\mathcal{L}=-\sum_{j=1}^{|\mathcal{V}|}\mathbbm{1}_{i=j}\cdot log(Q_{j})$.
4.2 Experimental Setup and Results
We experiment with training one layer of $\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}$ at a time. When training the layer that contains the edited MLP matrix $W^{\prime}_{V_{i}}$, we update all weights except $W^{\prime}_{V_{i}}$ (i.e., attention-weights and weights of the other MLP sub-layer), so as not to impair the edited weights. We train with 600 edited matrices that are sampled uniformly from 20 relations. We use 100 matrices from the same relations as a validation set. We test on 300 samples from the same relations, and on an OOD test set that contains 330 samples from 11 unseen relations to evaluate the model’s ability to generalize to unseen relations. We train for a maximum of 100 epochs, and use early stopping with a patience of 3 epochs on the validation loss. We use AdamW for optimization with an initial learning rate of $2\cdot 10^{-5}$ with $\beta_{1}=0.9,\beta_{2}=0.98$ and weight decay of $0.01$. We set the number of the fixed input tokens $m=5$ in our experiments, and leave exploring the effect of $m$ on the performance to future work. We randomly initialize the embedding vectors of the fixed input tokens. We evaluate based on the edited object accuracy (Meng et al., 2022), i.e., the accuracy of the model in generating the edited object $o^{\prime}_{i}$ based on $W^{\prime}_{V_{i}}$. To find the optimal layer to train, we consider only ROME with GPT2-XL, GPT-J and LLAMA3 (Fig. 3). Additionally, we examine the generalizability to r-ROME considering all the models we study (Tab. 1).
@@ Results.
The results in Fig. 3 show that the edited object can be generated with high accuracy (99% for the GPT-models and $97\%$ for LLAMA3 on CounterFact), when training a layer up to the layer containing the edited matrix. Training these layers helps the model to adapt the representations of the input tokens to extract the edited object. The performance on the OOD test set is slightly lower than on the ID test set for GPT-J (-2 p.p.) and LLAMA3 (-3 p.p.). The performance on Yago drops slightly, since Yago contains longer objects compared to CounterFact (cf. Tab. 15). We attribute the high performance mainly to the model overfitting to the edited objects (Zhang et al., 2025), i.e., the edited object having overly high probability after editing. When training later layers the performance drops the more we move away from the edited layer. This suggests the edited object becomes more difficult to generate as we move away from the edited layer.
Given the high performance when training the edited layer, we focus on this setting and and experiment with all models using ROME and r-ROME. We run each combination (editing method and model) with 5 random seeds. The results in Tab. 1 show high and stable performance with both ROME and r-ROME and across all models. For example, the in-domain accuracy is $88\%$ and the OOD accuracy $85\%$. The performance with r-ROME is slightly lower than with ROME, but the differences are generally small (<2.7<2.7 $p.$).
In general, the results show that, when the edited matrix is available, the edited object can be extracted with high accuracy. Our method provides direct information about the edit (the edited object $o^{\prime}$) with strong generalization, and can be combined with information about the relation (cf. App. D) to reconstruct the edited fact. We show that EditScope can generalize to other KEs in App. E.1
Figure 3: Accuracy of generating the edited object based on the edited matrix when training different layers of ROME-edited models (Left: CounterFact, Right: Yago).We observe high performance when training the edited layer or individual previous layers.
Method Model Acc. Std Acc. (OOD) Std (OOD)
ROME GPT2-XL 99.40 0.43 99.70 0.30
GPT-J-6B 97.60 1.86 94.42 1.51
META-LLAMA-3-8B 96.47 0.56 91.21 2.77
QWEN2.5-7B 91.20 2.06 87.45 2.73
r-ROME GPT2-XL 99.73 0.28 99.70 0.52
GPT-J-6B 96.50 2.86 95.91 3.37
META-LLAMA-3-8B 94.87 1.07 88.18 3.04
QWEN2.5-7B 88.53 1.71 85.45 4.00
Table 1: Accuracy of generating the edited object based on the edited matrix when training only the edited layer. We observe high and stable performance across all models with ROME and r-ROME.
5 Reversing Edits
To reverse edits, we exploit the fact that, to promote the edited object, it must be overly present in the edited matrix. We hypothesize that thereby particular rank-one approximations based on the highest singular values of a Singular Value Decomposition (SVD) of the edited matrix are similar to the rank-one update matrix of methods such as ROME or r-ROME. Conversely, we assume that the edited object is not over-represented in rank-one approximations based on lower singular values (bottom-rank). We introduce bottom-rank approximations derived from SVD in Sec. 5.1, conduct an analysis of our hypothesis in Sec. 4, and present our approach for reversing edits in Sec. 5.3.
5.1 Singular Value Decomposition and Bottom-Rank Approximations
We can further construct rank rkr-k approximations of $M$ by excluding the top (i.e., highest) $k$ singular values and their corresponding vectors from $U$ and $V$, and refer to these as bottom-rank approximations:
$\tilde{M}^{(r,k)}=\sum_{i=1}^{r}\mathbbm{1}_{\mathbf{i>k}}\tilde{M}^{(i)}$ (2)
5.2 Analysis of Rank-One Approximations
Figure 4: The maximum cosine similarity values between vectors of the update matrix $W_{N}$ and the rank-one approximation $\tilde{W^{\prime}}^{(k)}_{V}$.
Given that the update matrix $W_{N}$ makes the edited object quite prominent in the edited matrix, we hypothesize that some of the rank-one approximations of $W^{\prime}_{V}$ are similar to the rank-one update matrix $W_{N}$. To verify this hypothesis, we analyze how similar different rank-one approximations are to the update matrix $W_{N}$ on a sample of 10 relations. The row vectors of each rank-one matrix can have at most two directions. As proxy for similarity, we use the maximum cosine similarity value among the rows of $W_{N}$ and $\tilde{W^{\prime}}_{V}^{(k)}$ for different $k$ values. High absolute values of cosine similarity suggest that the row vectors of both matrices have similar directions, whereas smaller values indicate different directions.
@@
We show the results in Fig. 4 (extended by standard deviations in App. Tab. 16). The results show very high similarity (0.98) between the update matrix and the $k=1$ approximation for GPT2-XL. For larger $k$ values the similarity drops significantly. For GPT-J, the similarity with $k=1$ is lower (0.77), but we have a moderate similarity (0.45) with $k=2$. Here too, the similarity values drop when $k>2$. For LLAMA3, the values are much lower (0.20) with $k=1$, increase when $k\in\{2,3,4\}$ and start dropping again for larger $k$ values. For QWEN, we observe the same pattern as with GPT-J but with lower similarity values. This suggests that for GPT-models, the single rank-one approximation with the top singular value encodes the edit, whereas for LLAMA3, a combination of rank-one approximations from top singular values encodes the edit. In general, the results show that the rank-one approximations with $k=1$ come close to the update matrix in case of the GPT-models, whereas on LLAMA3 and QWEN the approximations have lower similarities to the update matrix.
$k$ GPT2-XL GPT-J-6B META-LLAMA-3-8B QWEN2.5-7B
Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$ Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$ Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$ Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$
0 0.00 100.00 0.32 100.00 0.97 100.00 0.65 100.00
1 87.10 7.42 32.26 60.65 5.48 95.48 31.94 62.90
2 88.39 4.84 72.90 6.77 28.39 66.45 51.94 42.58
3 90.32 2.90 76.77 5.81 44.84 50.00 53.55 40.00
4 90.32 1.94 75.81 6.13 60.97 28.39 53.87 37.10
5 91.29 1.94 77.42 3.23 66.77 20.32 56.77 34.19
6 91.29 1.94 77.10 2.90 67.74 18.71 58.71 30.97
7 90.97 1.94 77.42 2.58 71.29 13.87 59.68 30.32
8 91.29 1.94 77.74 2.58 73.23 11.94 59.68 30.32
9 92.58 1.94 78.06 2.58 75.16 9.68 60.65 29.03
10 93.87 1.94 78.06 2.58 76.77 9.03 62.90 27.42
11 94.52 1.29 78.06 2.58 76.77 8.71 62.58 26.45
12 94.19 1.29 79.03 2.58 79.35 7.10 62.90 26.13
13 93.23 0.97 79.35 2.26 79.35 6.77 62.90 26.13
14 93.55 0.97 80.00 2.26 78.71 6.77 62.58 25.16
15 93.87 0.97 78.71 2.26 80.00 6.45 62.58 24.52
Table 2: Reversal and editing accuracy with bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ for ROME. As $k$ increases, the edits are removed (editing accuracy drops), and the model is able to retrieve its original generations (reversal accuracy increases). Similar results for r-ROME and Yago are shown in App. Tab. 12 and Tab. 19 respectively.
5.3 Reversal
The results from the previous section suggest that the editing information might be localized at the first few rank-one approximations of $W^{\prime}_{V}$, and that the original object before editing might still be encoded in bottom-rank approximations of $W^{\prime}_{V}$. This observation encourages us to investigate replacing the edited matrix $W^{\prime}_{V}$ by its bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$. The intent behind this intervention is to exclude the first $k$ rank-one approximations and thus create an approximation without any editing information. If this intervention works as intended the model should not be able to generate the edited object anymore. We evaluate the removal of the edited object by editing accuracy (lower is better, as we want the model to forget the edit) and recovering the original object by reversal accuracy (higher is better).
Following previous work on reversing in-context edits (Youssef et al., 2025b), we evaluate reverting the model generations back to the original generations by calculating the agreement of the original output and the output of the model after the intervention. Editing and reversal accuracy are calculated as $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(\hat{y_{i}}=y_{i})$, where $\hat{y_{i}}$ is the reverse-edited output and $y_{i}$ is the original or edited output for edit $i$. For reversal accuracy, we approximate the model’s outputs using the next token prediction, following Du et al. (2024); Youssef et al. (2025b). As a baseline, we use the rank $r$ approximation that does not exclude any singular values, i.e., we set $k=0$. Here, we use 310 instances, uniformly sampled from 31 relations.